Jayden and Alexa were asked if 6C2 + 6C3 = 7C3 is true. Their calculations showed it was not true, but they both made mistakes in their calculations. Explain and correct their mistakes. I already corrected jaydens mistakes, but i cant figure out how to fix Alexa's. So please help. Alexa added the fractions by using the LCD (least common denominator).
Step 1 6! 6!
--- + ---
2!4! 3!3!
Step 2 6! 4*6!
---- + -----
4*3!3! 4*3!3!
Step 3 5*6!
------
4*3!3!
Step 4 7! 7*6!
---- = -----
3!4! 4*3!3!
Step 5 5*6! 7*6!
----- not equal -----
4*3!3! 4*3!3!

Answer by stanbon(75887)

6C2 + 6C3 = 7C3
(6*5)/(1*2) + (6*5*3/1*2*3) = (7*6*5/1*2*3)
-------
Multiply thru by 3! to get:
3*30 + 90 = 210
90 + 90 = 210
false
-----------------
Cheers,
Stan H.
-----------------